OpenStudy (anonymous):

8. Explain the meaning of the statement arccos(Sq.rt 3 )/2 = pi/6

OpenStudy (anonymous):

it means that \(\frac{\pi}{6}\)is the unique number between 0 and \(\pi\) whose cosine is \(\frac{\sqrt{3}}{2}\)

OpenStudy (anonymous):

i.e. \(\cos(\frac{\pi}{6})=\frac{\sqrt{3}}{2}\)